A 850-kg car moving at 12 m/s takes a turn around a circle with a radius of 22.0 m. Determine the acceleration and the net force acting upon the car.


1
Expert's answer
2021-10-26T17:14:52-0400

"m = 850\textsf{ kg}\\
v = 12\textsf{ m/s}\\
r = 22.0 \textsf{ m}\\
a =\ ?\\
F = \ ?"


"a =\dfrac{v²}r = \dfrac{12^2}{22} = 6.55\textsf{ m/s²}"


"F = \dfrac{mv^2}r = 850 × 6.55 = 5563.64\textsf{ N}"